OpenStudy (otherworldly):

Can someone please help me with this? Which statement is true about the product square root of 2(5 + square root of 8)? It is rational and equal to 7. It is rational and equal to 9. It is irrational and equal to 4 + square root of 10. It is irrational and equal to 4 + 5square root of 2.

OpenStudy (otherworldly):

@EclipsedStar

OpenStudy (danjs):

This what it looks like? \[\large \sqrt{2}*(5+\sqrt{8})\]

OpenStudy (otherworldly):

yes

OpenStudy (danjs):

do you recall how to distribute that to expand it?

OpenStudy (otherworldly):

i do not know how to distribute a square root

OpenStudy (danjs):

it is the same as if it was not a square root. a*(b + c) = a*b + a*c

OpenStudy (otherworldly):

oh okay so square root of 10 + square root of 16?

OpenStudy (danjs):

Distributing gives this to begin with \[\large \sqrt{2}*(5+\sqrt{8})=5*\sqrt{2}+\sqrt{8}*\sqrt{2}\]

OpenStudy (danjs):

You cant do anything to simplify the first term , but the second term can be simplified. rewrite the 8 as 2*4, and the square root of 4 is 2. \[\sqrt{8}* \sqrt{2}=\sqrt{4*2}*\sqrt{2}=\sqrt{4}\sqrt{2}*\sqrt{2}=2\sqrt{2}*\sqrt{2}=2*2=4\]

OpenStudy (danjs):

so overall the thing simplifies to \[\large \sqrt{2}*(5+\sqrt{8})=4+5\sqrt{2}\]

OpenStudy (otherworldly):

oh ok

OpenStudy (otherworldly):

can you tell me what does rational and irrational mean @DanJS ?

OpenStudy (otherworldly):

i just want to know for future problems

OpenStudy (danjs):

a number is rational if it can be written as a fraction of whole numbers a/b

OpenStudy (danjs):

if it cant be written as a/b ; then it is irrational

OpenStudy (otherworldly):

ok thank you so much

OpenStudy (danjs):

square root 2 is irrational... welcome